### v1

# SITransformer: Shared Information-Guided Transformer for Extreme Multimodal Summarization

Abstract.
Extreme Multimodal Summarization with Multimodal Output (XMSMO) becomes an attractive summarization approach by integrating various types of information to create extremely concise yet informative summaries for individual modalities. Existing methods overlook the issue that multimodal data often contains more topic irrelevant information, which can mislead the model into producing inaccurate summaries especially for extremely short ones. In this paper, we propose SITransformer, a Shared Information-guided Transformer for extreme multimodal summarization. It has a shared information guided pipeline which involves a cross-modal shared information extractor and a cross-modal interaction module. The extractor formulates semantically shared salient information from different modalities by devising a novel filtering process consisting of a differentiable top-k selector and a shared-information guided gating unit. As a result, the common, salient, and relevant contents across modalities are identified. Next, a transformer with cross-modal attentions is developed for intra- and inter-modality learning with the shared information guidance to produce the extreme summary. Comprehensive experiments demonstrate that SITransformer significantly enhances the summarization quality for both video and text summaries for XMSMO. Our code will be publicly available at https://github.com/SichengLeoLiu/MMAsia24-XMSMO.
summarization; multimodal learning; information denoising
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Video summarizationccs: Computing methodologies Information extraction
1. Introduction
In the digital age, information overload is a significant challenge due to the rapid production of data that surpasses human processing abilities. Summarization techniques aim to distill essential information quickly, aiding in managing this overload. Traditional text summarization and video summarization methods have been extensively studied (Kupiec et al., 1995; Liu and Lapata, 2019; Zhang et al., 2016). However, with the rise of multimodal content on the Internet, conventional unimodal summarization methods are increasingly inadequate. Hence, multimodal summarization with multimodal output (MSMO) (Zhu et al., 2018), which integrates information from multiple modalities, has emerged to address this gap. By incorporating diverse data sources such as videos and text documents, with advanced deep learning techniques (Vaswani et al., 2017; Hu et al., 2019), it enhances summarization quality and user experience. To further simplify and expedite information acquisition, extreme summarization was introduced into the MSMO task (Tang et al., 2023b, a), proposing eXtreme Multimodal Summarization with Multimodal Output (XMSMO). XMSMO condenses a video-document pair into a single frame and a single sentence. This extremely concise multimodal output allows users to quickly grasp the essence and decide whether to engage further.
Multimodal data contains rich information from various sources and formats to create more informative summary. However, the wealth of information can include noise and irrelevant details that may mislead or complicate the summarization process. For instance, as depicted in Figure 1, prolonged scenes of a host introducing news in a studio setting can be misleading if the model focuses solely on visual data. These scenes shown in red boxes do not reflect the actual news content and are typically irrelevant to the summary. Such challenges become more significant when generating extremely short summarization in the XMSMO task. Even minor errors can significantly reduce the quality of the summarization, underscoring the importance of accurately identifying topic-relevant information.
Figure 1. An example of irrelevant noises (in red box) and shared salient information (in green box).
To address this issue, we posit that shared information across different modalities is capable to provide critical insights for comprehending the topic, thereby facilitating better utilization of the relevant information from both shared and non-shared contents for the summary. Thus, this paper proposes an innovative multimodal summarization method called SITransformer for the video-document XMSMO task. Overall, SITransformer consists of a cross-modal share information extractor and a cross-modal interaction module for the effective exploitation and utilization of shared information guidance. Specifically, the extractor aims to extract the salient information shared across modalities by a differentiable top-k selection, which ranks and selects the most salient features across modalities. The shared information is then used to control the information gate units, which determine which portions of the source features should be retained and which should be discarded, thereby filtering out features irrelevant to the topic and summarization. The cross-modal interaction module then processes the extracted relevant information using visual and textual transformers with cross-modality attentions. This enables both unimodal and cross-modal information learning to produce the final summarization.
The main contributions of this paper are as follows:
  1. (1)
    We propose a novel summarization method, SITransformer, for the XMSMO task using cross-modal shared salient information.
  2. (2)
    We design an innovative cross-modal shared information extractor that includes a differentiable top-k selector and an information gating unit for systematically exploiting relevant multimodal information.
  3. (3)
    Comprehensive experiments on a large-scale multimodal dataset demonstrates the effectiveness of SITransformer in achieving state-of-the-art performance.
2. Related Work
Unimodal Summarization focuses on summarizing content from a single modality, such as text or video. In text summarization, extractive methods select key sentences from the original document (Liu, 2019). Abstractive methods instead generate new sentences, often using Transformer-based models like BART (Lewis et al., 2019). Hybrid methods, such as the Pointer-Generator Network (See et al., 2017; Tang et al., 2022), combine extraction and generation of content for the final summary. Video summarization involves selecting informative frames to represent the main content. Early studies relied on methods such as LSTMs to formulate temporal patterns (Zhang et al., 2016), while recent methods used Transformer-based methods to capture frame relationships (Zhao et al., 2022). These unimodal methods establish a critical foundation for multimodal summarization.
Multimodal Summarization aims to improve summarization quality and user experience by summarizing information from two or more modalities. Early studies typically utilized data from other modality as the auxiliary information to facilitate the summarization regarding the target modality (Liu et al., 2022; Zhu et al., 2023). MSMO was first introduced by (Zhu et al., 2018, 2020), where visual attention mechanisms were designed to identify the most salient images. Li et al. (2020) pioneered the incorporation of video into MSMO by utilizing a recurrent neural network (RNN) to capture temporal patterns, instead of extracting a summary from a set of individual images. Zhang et al. (2022) introduced a hierarchical cross-modality semantic correlation learning model (HCSCL) that employed a Faster R-CNN to detect visual objects. This method integrated object and text features as low-level information, and hierarchically fused sentences and scenes as high-level information. Such hierarchical fusion strategy has been used in many recent studies and proved to be an effective cross-modal information understanding method (Tang et al., 2023b; Xu et al., 2023).
Figure 2. The overall structure of the SITransformer, which consists of a multimodal embedding module, a cross-modal shared information extractor, a cross-modal interaction module and modality-specific extreme summarization decoders.
Extreme summarization condenses original data into extremely concise content. For instance, extreme video summarization aims to distill a video into a single frame, while extreme text summarization focuses on summarizing a document into one sentence. Tang et al. (2023b, a) introduced the XMSMO task, which summarizes a video-document pair into a single sentence and a single frame. Their hierarchical model integrated information across modalities at sentence-scene and video-document levels. By minimizing Wasserstein distances between selected frames/texts and original videos/documents respectively, the model improves summarization coherence. These studies offer a feasible approach for extreme summarization but are often affected by irrelevant information in rich multimodal data, resulting in less decent summarization quality.
3. Methodology
The overall pipeline of the proposed SITransformer is depicted in Figure 2. It consists of four core stages with the shared salient information guidance. Firstly, a multimodal embedding module, which consists of multimodal encoders, is used to embed the input video and document into latent features. Next, a cross-modal shared information extractor processes the latent features with a noise filtering process. A cross-modal interaction module further guides the summarization of each modality with the extracted shared information. Finally, an extreme summarization decoder creates the corresponding extreme summaries for video and document.
3.1. Problem Definition
The XMSMO task aims to summarize an input video-document pair into an extremely short one-sentence and one-cover-frame summary. Mathematically, each paired input can be denoted as $(\mathbf{T},\mathbf{V})$. We represent the text document as $\mathbf{T}=\{\mathbf{t}_{i}|i\in[1,n]\}$ and the video as $\mathbf{V}=\{\mathbf{v}_{j}|j\in[1,m]\}$. Here, $\mathbf{t}_{i}$ denotes the $i$-th word in the document $\mathbf{T}$, and $\mathbf{v}_{j}$ represents the $j$-th frame in the video $\mathbf{V}$. The $m$ and $f$ indicate the corresponding lengths, respectively. The summarization outputs are defined as a single sentence $\mathbf{O}_{T}=\{\mathbf{t}^{\prime}_{1},\ldots,\mathbf{t}^{\prime}_{o}\}\in\mathbf{T}$ and a cover-frame $\mathbf{O}_{V}=\{\mathbf{v}^{\prime}_{o}\}\in\mathbf{V}$ where $\mathbf{t}^{\prime}$ and $\mathbf{v}^{\prime}$ are created based on the input document and video.
3.2. Multimodal Embedding Module
To achieve a basic alignment between visual and textual modalities, we apply the widely used multimodal embedding model, CLIP (Radford et al., 2021), as our latent feature encoder. The CLIP’s multimodal latent space enables the visual and textual contents with related semantics to share similar embeddings, which is able to facilitate the cross-modal shared information exploration.
Mathematically, for the text input $\mathbf{T}$ and video input $\mathbf{V}$,their word and frame-level low-level features, denoted as $\mathbf{X}_{T}^{L}$ and $\mathbf{X}_{V}^{L}$, respectively, can be obtained as:
(1) $\mathbf{X}_{T}^{L}$ $=\mathrm{CLIP}(\mathbf{T})=\{\mathbf{x}_{t,i}^{l}|i\in[1,n]\}\in\mathbb{R}^{n\times d},$
$\mathbf{X}_{V}^{L}$ $=\mathrm{CLIP}(\mathbf{V})=\{\mathbf{x}_{v,j}^{l}|j\in[1,m]\}\in\mathbb{R}^{m\times d},$
where $d$ is the latent dimension size, $n$ denotes the number of words, and $m$ denotes the number of video frames.
In addition to the low-level features (word and frame-level), high-level features (video and document-level), $\mathbf{X}_{T}^{H}$ and $\mathbf{X}_{V}^{H}$ are further considered for both document and video by using a mean pooling method across all words and frames, respectively. The video and document-level features can be obtained as:
(2) $\mathbf{X}_{T}^{H}$ $=\mathrm{pool}(\mathbf{X}_{T}^{L})\in\mathbb{R}^{d},$
$\mathbf{X}_{V}^{H}$ $=\mathrm{pool}(\mathbf{X}_{V}^{L})\in\mathbb{R}^{d}.$
3.3. Cross-Modal Shared Information Extractor
The low and high-level features contain rich information relevant to the major topic. However, they often include a significant amount of irrelevant information, making precise summarization challenging. To reduce such negative impact, a cross-modal shared information extractor with noise filtering mechanism is devised. The core idea is to leverage shared salient information, presenting in both modalities, to regulate information gates and filter out irrelevant information from the shared and non-shared feature embeddings accordingly.
Figure 3. Noise Filter with Differentiable Top-k (NFDT).
The noise filtering mechanism is designed with a Differentiable Top-k (NFDT) (Cordonnier et al., 2021) strategy as depicted in Figure 3. Specifically, linear layers are first adopted to obtain salient information scores for each word and frame:
(3) $\mathbf{s}_{T}$ $=\mathbf{W}_{1}\mathbf{X}_{T}^{L}+\mathbf{b}_{1}\in\mathbb{R}^{n},$
$\mathbf{s}_{V}$ $=\mathbf{W}_{2}\mathbf{X}_{V}^{L}+\mathbf{b}_{2}\in\mathbb{R}^{m},$
where $\mathbf{S}_{T}=\{s_{t,i}|i\in[1,n]\}$ and $\mathbf{S}_{V}=\{s_{v,m}|i\in[1,f]\}$ are the scores for each word and frame; $\mathbf{W}_{1}$, $\mathbf{W}_{2}$, $\mathbf{b}_{1}$ and $\mathbf{b}_{2}$ are learnable weight and bias, respectively.
Next, Gumbel softmax functions (Jang et al., 2016) are utilized to standardize the score between 0 and 1. Compared to the regular softmax function, the Gumbel softmax function is differentiable by integrating noise sampled from Gumbel distribution. Mathematically, we first sample noise $g$ by:
(4) $g=-\mathrm{log}(-\mathrm{log}(u))\in\mathbb{R},$
where $u$ is a random variable uniformly distributed on $[0,1]$. Gumbel softmax is then calculated as follows:
(5) $\text{score}_{t,i}$ $=\frac{\mathrm{exp}((s_{t,i}+g_{i})/\tau)}{\sum_{i=1}^{n}\mathrm{exp}((s_{t,i}+g_{i})/\tau)}\in\mathbb{R},$
$\text{score}_{v,j}$ $=\frac{\mathrm{exp}((s_{v,j}+g_{j})/\tau)}{\sum_{j=1}^{m}\mathrm{exp}((s_{v,j}+g_{j})/\tau)}\in\mathbb{R},$
where the $\text{score}_{t,i}$ and $\text{score}_{v,j}$ are the final scores for the $i$-th word and the $j$-th frame, respectively; $\tau$ is a temperature parameter that controls the smoothness.
We can obtain the score list $\mathbf{S}_{T^{\prime}}=\{score_{t,i}|i\in[1,n]\}$ and $\mathbf{S}_{V^{\prime}}=\{score_{v,j}|j\in[1,m]\}$ for all individual texts and frames. The top-k features with highest estimated scores are selected to construct the potential multimodal shared salient information $\mathbf{X}_{S}$ as:
(6) $\mathbf{X}_{S}$ $=\mathrm{Top}_{k}(\mathbf{S}_{T^{\prime}})\cup\mathrm{Top}_{k}(\mathbf{S}_{V^{\prime}})$
=$=\{\mathbf{x}^{l}_{t,n_{1}},\ldots,\mathbf{x}^{l}_{t,n_{k}}\}\cup\{\mathbf{x}^{l}_{v,m},\ldots,\mathbf{x}^{l}_{v,m_{k}}\}\in\mathbb{R}^{2|k|\times d},$
where $k$ is a hyper-parameter that controls the amount of selected multimodal shared salient information.
Overall, the multimodal shared salient information $\mathbf{X}_{S}$ contains $k$ image and word features simultaneously. It will be used to concatenate with word and frame features respectively to provide shared information guidance to the irrelevant information filtering process. The shared information is first pooled as:
(7) $\mathbf{x}_{S}=\mathrm{pool}(\mathbf{X}_{S})\in\mathbb{R}^{d}.$
The pooled information is then used to control an novel information gating unit. Specifically, if the gate vector is 0, the feature is considered irrelevant and discarded; otherwise, it is retained. The gate vector is defined as:
(8) $\mathbf{G}_{T}$ $=\mathrm{sigmoid}(\mathbf{W}_{3}(\mathbf{x}_{S}\oplus\mathbf{X}_{T}^{L})+\mathbf{b}_{3})\in\mathbb{R}^{n\times d},$
$\mathbf{G}_{V}$ $=\mathrm{sigmoid}(\mathbf{W}_{4}(\mathbf{x}_{S}\oplus\mathbf{X}_{V}^{L})+\mathbf{b}_{4})\in\mathbb{R}^{m\times d},$
where $\mathbf{G}_{T}$ and $\mathbf{G}_{V}$ are the gate matrices for text and frames, respectively. $\mathbf{W}_{3}$, $\mathbf{W}_{4}$, $\mathbf{b}_{3}$ and $\mathbf{b}_{4}$ are learnable parameters. Here, $(\mathbf{x}_{S}\oplus\mathbf{X}_{T}^{L})$ and $(\mathbf{x}_{S}\oplus\mathbf{X}_{V}^{L})$ represent the broadcasting of $\mathbf{x}_{S}$ and its concatenation to the beginning of each row of $\mathbf{X}_{T}^{L}$ and $\mathbf{X}_{V}^{L}$, respectively.
Finally, we filter out irrelevant information with the shared salient information guided gate by dot multiplying the original text and frame features with the corresponding gate vector.
(9) $\mathbf{X}_{T_{clean}}^{L}$ $=\mathbf{X}_{T}^{L}\cdot\mathbf{G}_{T}\in\mathbb{R}^{n\times d},$
$\mathbf{X}_{V_{clean}}^{L}$ $=\mathbf{X}_{V}^{L}\cdot\mathbf{G}_{V}\in\mathbb{R}^{m\times d},$
where $\mathbf{X}_{T_{clean}}^{L}$ and $\mathbf{X}_{V_{clean}}^{L}$ are the noise-free word and frame features, respectively.
3.4. Multi-Level Unimodal Representations
Based on the salient low-level features, unimodal representations can be finally obtained jointly with the high-level features. For the input document, document- and word-level features are concatenated first. Next, a Transformer is utilized to introduce a video-level understanding to all frame-level features:
(10) $\mathbf{X}_{T}^{\prime}$ $=\mathbf{X}_{T_{clean}}^{L}\oplus\mathbf{X}_{T}^{H}\in\mathbb{R}^{(n+1)\times d},$
$\mathbf{X}_{T}$ $=\mathrm{Transformer}(\mathbf{X}_{T}^{\prime})\in\mathbb{R}^{n\times d}.$
Note that we discard the high-level information token in the output as it has already exchanged information with the low-level information. Similarly, we can process the frame-level features as:
(11) $\mathbf{X}_{V}^{\prime}$ $=\mathbf{X}_{V_{clean}}^{L}\oplus\mathbf{X}_{V}^{H}\in\mathbb{R}^{(m+1)\times d},$
$\mathbf{X}_{V}$ $=\mathrm{Transformer}(\mathbf{X}_{V}^{\prime})\in\mathbb{R}^{m\times d}.$
3.5. Cross-Modal Interaction Module
To facilitate the multimodal learning and explore the cross-modal interaction, a cross-attention based mechanism is adopted. We formulate the attention scores from (1) video to document and (2) document to video as:
(12) $\mathbf{A}_{V2T}$ $=\mathrm{Attention}(\mathbf{Q}=\mathbf{X}_{V},\mathbf{K}=\mathbf{X}_{T},\mathbf{V}=\mathbf{X}_{T})\in\mathbb{R}^{m\times d},$
$\mathbf{A}_{T2V}$ $=\mathrm{Attention}(\mathbf{Q}=\mathbf{X}_{T},\mathbf{K}=\mathbf{X}_{V},\mathbf{V}=\mathbf{X}_{V})\in\mathbb{R}^{n\times d}.$
This mechanism enables effective information interaction between video and document, enhancing multimodal understanding. Next, a multimodal embedding $\mathbf{X}_{O}$ can be obtained by:
(13) $\mathbf{X}_{O}=\mathbf{W}_{5}(\mathbf{A}_{V2T}\oplus\mathbf{A}_{T2V})+\mathbf{b}_{5}\in\mathbb{R}^{(m+n)\times d}.$
To produce summaries for each modality with the multimodal guidance, we leverage a Transformer to inject the guidance:
(14) $\mathbf{X}_{TO}=\mathrm{Transformer}(\mathbf{X}_{O}\oplus\mathbf{X}_{T})\in\mathbb{R}^{n\times d},$
$\mathbf{X}_{VO}=\mathrm{Transformer}(\mathbf{X}_{O}\oplus\mathbf{X}_{V})\in\mathbb{R}^{m\times d},$
where the $\mathbf{X}_{TO}$ and $\mathbf{X}_{VO}$ are the multimodal aware word features and frame features, respectively. Note that we discard the multimodal relevant tokens for the outputs.
3.6. Extreme Summarization Decoder
Based on the multimodal features $\mathbf{X}_{TO}$ and $\mathbf{X}_{VO}$, the one-sentence-summary and cover-frame are created by the corresponding extreme summarization decoders. For text summary, a linear layer is used to compute the scores for the words in the document. Top-$k^{\prime}$ words are selected from the input document for the generation of the one-sentence-summary $\mathbf{O}_{T}$:
(15) $\mathbf{O}_{T}=\mathrm{Top}_{k^{\prime}}(\mathrm{Softmax}(\mathrm{Linear}(\mathbf{X}_{TO}))),$
where $k^{\prime}$ is the maximum number of words in summary. The selected $k^{\prime}$ words are ranked by their scores from the word-wise linear layer with the softmax activation, allowing the summary sentence to be constructed by ordering these words accordingly.
The cover frame selection can be regarded as a classification task. We use a linear layer with softmax activation to score the frames and identify the one $\mathbf{O}_{V}$ with the maximum score as the cover frame:
(16) $\mathbf{O}_{V}=\mathop{\mathrm{argmax}}\limits_{j}(\mathrm{Softmax}(\mathrm{Linear}(\mathbf{X}_{VO}))).$
3.7. Loss and Optimization
The proposed method is trained in an unsupervised manner. We employ the Wasserstein distance to assess the fidelity of generated summaries, a method well-established in prior research (Arjovsky et al., 2017; Tang et al., 2023b, a). This metric helps measure summarization quality as distributional differences. The calculation of Wasserstein distance between two distributions $p$ and $q$ is:
(17) $\begin{gathered}W\left(p,q\right)=\underset{\mathbf{\Gamma}\left(p,q\right)}{\operatorname{argmin}}\sum_{i,j}\gamma_{ij}\left(p_{i},q_{j}\right)c_{ij},\\ \text{ s.t. }\sum_{j=1}^{d}\gamma_{ij}\left(p_{i},q_{j}\right)=p_{i},\sum_{i=1}^{d}\gamma_{ij}\left(p_{i},q_{j}\right)=q_{j},\\ \gamma_{ij}\left(p_{i},q_{j}\right)\geq 0,\forall i,j.\end{gathered}$
where $c_{i,j}$ is the Euclidean distance based cost between $p$ and $q$, $d$ is the number of features. $\gamma_{ij}$ represents the mass moved from $p_{i}$ to $q_{j}$, where $p_{i}$ and $q_{j}$ represent the discrete points in the feature spaces associated with the distributions $p$ and $q$, respectively.
Following Tang et al. (2023b), we adopt three types of Wasserstein distance losses to represent the optimal transport from predicted distribution to desired distribution as follows:
(18) $\mathcal{L}_{T}=W(p_{T},q_{T}),\mathcal{L}_{V}=W(p_{V},q_{V}),\mathcal{L}_{O}=W(p_{O},q_{O}).$
$\mathcal{L}_{T}$ calculates the distributional distance between the one-sentence-summary and the input document, where $p_{t}$ and $q_{t}$ are probability mass functions of document feature $\mathbf{X}^{H}_{T}$ and summary feature $\mathbf{X}_{TO}$. Similarly, $\mathcal{L}_{V}$ represents the distance between cover-frame summary and the input video probability mass function. $\mathcal{L}_{O}$ further ensures the output summary is consistent between sentences and frames.
In addition to summary accuracy, we use Syntactic Log-Odds Ratio (SLOR) (Kann et al., 2018) to ensure the fluency of the textual summary as:
(19) $\mathcal{L}_{f}=\frac{1}{|\mathbf{\mathbf{O}_{T}}|}(\mathrm{ln}(P_{LM}(\mathbf{O}_{T}))-\mathrm{ln}(P_{U}(\mathbf{\mathbf{O}_{T}}))),$
where $P_{LM}(\mathbf{O}_{T})$ is the probability of the one-sentence assigned by a large language model (GPT-2 in practice). $P_{U}(\mathbf{O}_{T})$ is the unigram probability, which can be calculated as:
(20) $P_{U}(\mathbf{O}_{T})=\prod_{t\in\mathbf{O}_{T}}P(t).$
To this end, the overall loss can be calculated as:
(21) $\mathcal{L}=\lambda_{T}\mathcal{L}_{T}+\lambda_{V}\mathcal{L}_{V}+\lambda_{O}\mathcal{L}_{O}+\lambda_{f}\mathcal{L}_{f},$
where $\lambda_{T}$, $\lambda_{V}$, $\lambda_{S}$ and $\lambda_{f}$ are hyper-parameters that used to control the importance of each loss term.
4. Experiments & Discussions
4.1. Experimental Settings
Dataset. We adopted the multimodal news dataset proposed in Tang et al. (2023b). It contains 4,891 BBC News video-document pairs. The average video duration is 345.5 seconds and the average length of documents is 498 words. We extracted one frame every 120 frames, and a total of 60 frames were extracted as the video inputs. All video frames were resized to 640*360.
Methods for Comparison. We compared our SITransformer with the state-of-the-art unimodal extreme summarization methods (See et al., 2017; Qi et al., 2020; Zhang et al., 2020; Apostolidis et al., 2021, 2022), MSMO (Li et al., 2020; Radford et al., 2021) and XMSMO methods (Tang et al., 2023a, b).
Evaluation Metrics. Following existing practice (Tang et al., 2023b), for the quantitative evaluation of video summarization, intersection over union (IoU) (Sharghi et al., 2017) ad frame accuracy (FA) (Messaoud et al., 2021) are adapted. In addition, ROUGE (Lin, 2004) metric is adopted for the text summarization.
Implementation Details. We used AdamW (Loshchilov and Hutter, 2017) as the optimizer with a learning rate of 0.001 and a 12-layer Transformer with 16 attention heads for both unimodal data processing and cross-modal interaction. The hidden size of the Transformer was set to 512. The temperature parameter $\tau$ used in NFDT was set to 0.5.
4.2. Quantitative Analysis
Methods Visual $\uparrow$ Textual $\uparrow$
FA IoU R-1 R-2 R-L
PG (See et al., 2017) - - 2.43 0.08 2.25
ProphetNet (Qi et al., 2020) - - 3.77 0.09 3.56
PEGASUS-XSUM (Zhang et al., 2020) - - 4.36 0.12 4.00
CA-SUM (Apostolidis et al., 2022) 0.57 0.69 - - -
ARL (Apostolidis et al., 2021) 0.59 0.68 - - -
VMSMO (Li et al., 2020) 0.57 0.69 divg divg divg
CLIP (Radford et al., 2021) 0.58 0.68 3.35 0.05 3.14
TLDW (Tang et al., 2023b) 0.57 0.68 4.64 0.07 4.33
TopicCAT (Tang et al., 2023a) 0.60 0.69 4.54 0.08 4.21
Ours 0.64 0.79 6.37 0.09 6.11
w/o Shared info sel. 0.58 0.7 5.24 0.06 5.10
w/o Gumbel-softmax 0.61 0.73 5.83 0.07 5.38
w/o Shared info gate 0.59 0.73 5.02 0.01 4.54
w/ Random selector 0.59 0.7 4.37 0.02 3.93
w/ NFCS 0.61 0.65 5.67 0.06 5.23
Table 1. Quantitative evaluation on the multimodal BBC news dataset (Tang et al., 2023b).
Table 1 lists the quantitative results. Overall, our SITransformer outperforms the state-of-the-art methods regarding both extreme visual and textual summarization metrics. For the visual summarization, SITransformer achieves a FA of 0.64 and an IoU of 0.79. The increase in FA and IoU highlights the the SITransformer’s efficacy in identifying the most informative frame. Moreover, SITransformer achieves 6.37, 0.09, 6.11 on ROUGE-1, ROUGE-2 and ROUGE-L, respectively. The improvements in ROUGE metrics indicate the SITransformer’s enhanced ability to capture keywords and create more fluent sentences. Compared to unimodal extreme summarization methods, SITransformer shows better performance in both text and video summarization, highlighting the advantages of using multimodal data. Moreover, our approach outperforms all existing (X)MSMO methods, demonstrating the effectiveness of leveraging shared salient information for the summarization process.
4.3. Qualitative Analysis
Figure 4. Qualitative examples by our NFDT and the state-of-the-art methods: TopicCAT (Tang et al., 2023a), TLDW (Tang et al., 2023b) and CLIP (Radford et al., 2021).
We compare a number of summaries produced by our method, TopicCAT (Tang et al., 2023a), CLIP (Radford et al., 2021), TLDW (Tang et al., 2023b) and the ground truth in Figure 4. From the visual perspective, the cover frame selected by SITransformer is more consistent with the ground truth, while the selection of baselines will be affected greatly by irrelevant information. In the three examples, SITransformer selects cover frames with semantics similar to the ground truth, such as a Koala, while the baseline struggles to distinguish key information from irrelevant details and mistakenly selects irrelevant frames, like the news narrator, as cover frames.
As for the text summary, SITransformer is able to better capture the essence of the documents corresponding to the ground truths. However, all methods could be struggling in language fluency and grammatical errors, leading to room of improvement. Such problems are common for extractive summarization methods as they prioritize capturing keywords over language fluency(Tang et al., 2023a). Consequently, extractive summaries emphasize topic coverage, leading to a trade-off between fluency and coverage.
4.4. Ablation Study
We conducted ablation studies to verify the effectiveness of each proposed components. The experimental results are listed in Table 1. Firstly, we removed the shared information selection process in obtaining the multimodal shared salient information. The significantly degraded performance indicates the importance of extracting shared salient information to guide the summarization. We also assess the impacts of Gumbel-softmax by replacing with standard softmax, and evaluate shared information guided gate design by its removal in our proposed NFDT. The results demonstrate their positive contributions to enhancing the extraction of shared salient information. Lastly, to verify the effectiveness of the proposed noise filter design, we replace our NFDT noise filter with two different types of filters, namely random feature selector and noise filter with cosine similarity (NFCS). The former filters shared salient features by random selection while the later applies cosine similarity based ranking for salient information identification. Compared to our best model, both visual and textual summary metrics degrade, demonstrating the performance of our proposed novel noise filter.
Figure 5. Impacts of shared salient information amount in percentage on quantitative metrics
4.5. Impacts of Shared Information Amount
The amount of shared salient information to extract is an important hyper-parameters. We visualize its impacts during the multimodal shared information extraction process in Figure  5. It shows that retaining 50% of shared information leads to better performance in terms of most metrics except IoU. Selecting either an excessive or insufficient number of features can adversely affect performance. Choosing too many shared salient features may result in the inclusion of undesirable, irrelevant information, leading to the loss of valuable data. Conversely, using too few shared salient features may fail to adequately control the noise filter, also causing performance degradation.
4.6. Limitations and Future Work
SITransformer can produce sub-optimal text summarization for some scenarios, exhibiting grammatical errors and structural inconsistencies. One potential solution is to leverage large language models (LLMs) such as BLIP2 (Li et al., 2023) and MiniGPT5 (Zheng et al., 2023) to further refine the fluency and grammar accuracy of the results. Furthermore, future work could explore methods that generate cover frames based on user preferences and customization, using diffusion methods (Rombach et al., 2022), which further improve the flexibility of XMSMO.
5. Conclusion
This paper presents SITransformer for XMSMO, which leverages cross-modal shared information to guide the summarization process. A cross-modal shared information extractor is devised to identify shared salient information and a transformer based cross-modal interaction module facilitates the cross-modality learning. Our extensive experiments have demonstrated the effectiveness of our proposed method.